# The physical basis of teleological evolutions

### v4

@@ 2.2 Teleological character of natural evolutions – a possible physical basis
The second part of the work examines the possibility that the time-symmetric quantum causal loops underlying the teleological character of quantum algorithms are also the missing physical basis of the teleological character of the natural evolutions for which the teleological notion was conceived.
@@
This, of course, immediately raises the problem of explaining the existence of quantum features – indeed the quantum speedup – in evolutions that can be the size of the Earth and have a duration of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Naturally, at the same time we should adopt quantum cosmology $\left[13,14\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
@@
An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[15-19\right]$, the “Fine-tuned Universe” version of it. This version of the Principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”. Let us assume, for simplicity, that this correlation today is one-to-one. The significance of this principle lies in the fact that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
Under a quantum cosmological interpretation of it, the correlation addressed by the Anthropic Principle becomes quantum. We should think of a value of the fundamental constants observable and a value of life observable. At the beginning, there should be a quantum superposition (or mixture) of all the possible universes, each with its value of the fundamental constants. An initial measurement of the value of the fundamental constants observable would select one value at random. The subsequent unitary evolution of the universe would produce, at the present time, the corresponding value of life, which would then be measured. In Subsection 5.2., after the formalization of the notion of the time-symmetric quantum causal loops, we will provide a formalized toy model of the building up of the correlation in question on the part of the universe’s evolution.
By the way, since the value of the fundamental constants never changes along the universe’s evolution, in the universe we are in we can do without the initial measurement (when there were no observers). Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life. The measurement of either observable today would select the eigenvalues of both maximally entangled observables. Everything is as if the value of the fundamental constants was selected at the beginning of time.
Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants). We have seen that this unitary evolution can always occur with quantum speedup – there is the room for it. Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[12\right]$. It should consequently be teleological in character.
In the universe we are in, the search for one-to-one correlation between the value of the fundamental constants and the value of life becomes search for sentient life. It is thus the universe’s evolution toward sentient life that occurs with quantum speedup and is consequently teleological in character.
@@
In principle, the evolution of life on Earth can be obtained from that of the universe toward sentient life by tracing out all the rest. Evolution times would be identical and the amount of the resources employed by the evolution of life on Earth could not exceed those employed by its image in the evolution of the universe that comprises them. In the end, the evolution of life on Earth, even if decohered, would occur with quantum speedup and would thus be teleological in character. This should also hold for the evolutions of species, etc..
Of course, all the above is puzzling.
A physical basis to the dismissed notion of teleological evolution would be a paradigm shift for the natural sciences. As is well known, empirically, the teleological notion better fits with the functional structure and the evolutions of the living. If it has been discarded, it is only because of the assumed absence of a physical basis. For the time being, the fact that quantum algorithms are certainly teleological evolutions, and that, under a quantum cosmological interpretation of the Anthropic Principle, natural evolutions would be teleological in character too, may shake the current belief that such a physical basis cannot exist.
That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the usual strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence. Things become quite different in the case of the evolutions of the living. The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling. It might also go along with ancient beliefs, but of a kind deliberately ignored by science. Once science was accused of heresy for denying a physical basis to beliefs of the kind in question. We could be facing the reverse problem now: proposing a physical basis to the notion of teleological evolution could well be seen as a scientific heresy.
@@ 3 Positioning the work
We try and position the present work. Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Besides the retrocausal approach to the quantum computational speedup $\left[1-5\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
The present approach is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of being open to the violation of the ordinary notion of causality, not in common the way of looking at the quantum description of quantum algorithms. For example, in $\left[27\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time. So far so good but, consequently, they are excluded (by the approach in question) from the phenomena that can accommodate causality violations. Of course, for us, this means putting too much faith in their current quantum description. As we have seen, with their quantum computational speedups, quantum algorithms should be a most clear example of violation of one-way causality – what obviously implies changing their current quantum description.
@@
One is the local explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. Here, still locally, it would tell the other observable the state to assume in view of its future measurement. This retrocausal, local, explanation of quantum nonlocality is similar to ours. The difference is that sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics.
@@
Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[13\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating from before Deutsch’s 1985 discovery, the two principles in question make no reference to the quantum speedup of the universe’s evolution and, a-fortiori, its consequent teleological character.
4 The time-symmetric quantum causal loops – formalization
We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[6\right]$ quantum search algorithm. As anticipated in the Rationale, this immediately generalizes first to any number of drawers, then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. The formalization consists of the following steps:
@@ 4 The time-symmetric quantum causal loops – formalization
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We call this extended quantum description the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.

### v6

@@ 1.2 The possible teleological character of the evolutions of the living
The teleological character of quantum algorithms naturally raises the question of whether the quantum superposition of causal loops that underlie it can also be the missing physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived. Having established the teleological character of quantum algorithms, probing this possibility would even seem to be a duty of the physical science to the natural sciences.
@@
One difficulty, of course, would be to explain the presence of quantum features – the quantum speedup – in the evolutions of life that can be Earth-sized and last billions of years. However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels opens up another possibility: that they scale down from the largest and longest existing evolution, that of the universe. Of course, at the same time, we should adopt quantum cosmology $\left[12,13\right]$, i.e., the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
@@
Interestingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together. The Fine-tuned Universe version of this principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today (in particular: “presence of sentient life” or “absence of life”). The reason for this correlation is that the slightest variation in the value that the fundamental constants have in the universe we are in would produce a radically different universe, most likely incapable of hosting life. Let us assume, for simplicity, that this correlation is one-to-one.
The quantum interpretation of the Anthropic Principle produces a formal identity between the building up, by the unitary evolution of the universe under the fundamental physical laws and constants, of the quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today on the one side and, on the other, the building up, by a quantum search algorithm, of the quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end. In other words, the unitary evolution of the universe toward the value of life, formally, turns out to be a quantum search algorithm. These quantum algorithms yield a quantum speedup even if the search for the correlation is random $\left[11\right]$. It is indeed reasonable (and conservative) to assume that the search for the correlation between the value of the fundamental constants since the beginning of time and the value of life today by the universe’s evolution – the search for life in the universe we are in – is random.
@@
According to this formal identity, the evolution of the universe toward the value of life should occur with quantum speedup and would consequently be teleological in character. In the universe we are in, the same should hold for the evolution of the universe toward life. It is then easy to show that this teleological character should be inherited by the evolution (and the evolutions) of life on Earth, derivable from that of the universe by tracing out everything else. In fact, evolution times should be the same for the evolution of life on Earth and its image in the evolution of the universe and the resources employed by the former evolution cannot exceed those involved by the latter that comprises them.
Of course, that the universe evolves in a unitary quantum way, although consistent with the theory of decoherence, for which quantum processes decohere by losing information in their environment, is just an assumption. It should instead be firm the derivation from it of the teleological character of the evolutions of the living. In fact, the Fine-tuned Universe version of the Anthropic Principle is universally accepted and the quantum cosmological interpretation of it is straightforward.
Summing up, we would conclude with the conditional statement that, if quantum cosmology were right, the evolutions of the living would be teleological in character.
Let us also observe that this would be a paradoxical possibility. For example a way of making teleological a Darwinian evolution could be to assume that those random DNA mutations that, in the future, will turn out to survive because more suited to the environment, because of that, would become more probable. By the way, this would be exactly as paradoxical as the causal loop explanation of the quantum speedup. The only difference is that the quantum computational speedup belongs to a quantum world that is notoriously the place of queer phenomena. A teleological Darwinism would instead concern the macroscopic world we see around us. In effect, evolutions attracted by their future would be a radically different way of viewing physical reality.
In the next Section 2, we try to position the present work within the literature that challenges the still dominant notion of causal evolution. In Section 3, we review and further develop the quantum causal loop description of quantum algorithms. Section 4 is the (of course, speculative) transposition of the teleological description of quantum algorithms to natural evolutions. Section 5 is on the prospects for further research. Section 6 gives the conclusions.
@@ 2 Positioning the work
Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-27\right]$.
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the way of looking at quantum algorithms. For example, in $\left[26\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time. So far so good but, consequently, quantum algorithms are excluded from the phenomena that can accommodate causality violations. Of course, for us, this means putting too much faith in their current quantum description. To us, with their quantum computational speedups, quantum algorithms are the seminal example of causality violation.
@@
One is the celebrated EPR paper $\left[28\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description is incomplete. More specifically, they claimed that the quantum description of quantum nonlocality is incomplete because it mathematically predicts something that it does not physically explain, namely the “spooky action at a distance”. As we will see, time-symmetrizing the ordinary quantum description of quantum nonlocality, as we did with that of quantum algorithms, completes it. It generates a quantum superposition of causal loops that play the role of the hidden variables those authors envisaged to complete the quantum description.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
This retrocausal explanation of quantum nonlocality is similar to ours as it also involves a causal loop. The difference is in the character of the causal loops. Sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics but is instrumental to finding a local explanation of quantum nonlocality. As already said, the causal loops originated by time-symmetrizing the ordinary quantum description are instead imposed by quantum mechanics. Therefore, first, they must be there independently of the fact of explaining quantum non-locality locally; second, as long as quantum mechanics is consistent, they should be exempt from the paradoxes that plague the usual kind.
@@
Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating to before Deutsch’s 1985 discovery, of course the two principles in question make no reference to a quantum speedup of the universe’s evolution and its consequent teleological character.
3 Completing the quantum description of quantum algorithms
This section is divided into two parts. The first begins with a constructive critique of the ordinary quantum description of Grover algorithm and ends with a generalization of the results obtained. We show (i) why the quantum description in question, being causal in character, must be considered incomplete, (ii) how it should be completed, and (iii) how both things extend to a very general class of quantum correlation processes. The second part formalizes the completion procedure consisting in a time-symmetrization of the ordinary quantum description.
3.1 Updating the EPR critique of the quantum description
Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[28\right]$. As is well known, Einstein in particular, argued that the ordinary quantum description of quantum nonlocality is incomplete because it does not physically explain the “spooky action at a distance” that it mathematically predicts. They also stressed the need to devise ”hidden variables” that would complete the quantum description.
The present critique of the quantum description could be considered an update of that of Einstein and the others because:
(i) it shows that Einstein et al.’s assertion that the quantum description of quantum nonlocality is incomplete can be extended to quantum algorithms, of course discovered exactly half a century after the publication of the EPR paper;
(ii) the reason for the incompleteness of the quantum description turns out to be the same in both cases;
(iii) in the quantum computational context, it turns out to be evident that the quantum description should be completed by time-symmetrizing it: completing in the same way the quantum description of quantum nonlocality, locally (physically) describes the spooky action at a distance.
First, we develop our critique of the quantum description on that of the four drawers instance of Grover quantum search algorithm $\left[1\right]$. For convenience, we summarize the part of this critique anticipated in the introduction.
In its simple instance, the problem solved by Grover algorithm is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. In the quantum case – by the optimal quantum search algorithm of Grover – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case.
It is natural to ask ourselves how it is possible that, in the quantum world, a ball hidden in one of four drawers can always be located by opening a single drawer. Of course, if one goes into the details of the quantum algorithm, one sees the ingenious interplays between quantum state superpositions and interferences that lead to producing the location of the ball by opening just one drawer. However, these interplays are naturally different from quantum algorithm to quantum algorithm and, as far as we know, in mainstream quantum information there is no known way of unifying a quantitative reason for which they yield a quantum speedup. A unified, quantitative, reason is provided in the retrocausal approach to the quantum speedup of ours and coauthors but, as things are now, this unconventional approach never took off within the scientific community.
In the next subsection we address the unification problem. We will see that a unifying quantitative explanation of the quantum speedups becomes quite evident once the quantum descriptions of quantum algorithms is complete. It is a unifying explanation since it is the same for all quantum algorithm. Since deriving it requires an unconventional way of reasoning, open of course to retrocausality, we will give first the rationale then, in subsection 3.2., the formalizations.
3.1.1 Grover quantum search algorithm as a quantum superposition of causal loops
We have already shown in the Introduction that the sheer existence of the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps). In the optimal quantum case, the number of computations steps is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question.
Always in the introduction, we wandered what was the impact of the above fact on the current quantum description of the quantum algorithm, noticing that one should take a quantum description not limited to the process of solving the problem but extended to the upstream process of setting the problem.
Our ordinary quantum description of the quantum algorithm therefore became:
An initial measurement in a quantum superposition (or mixture) of all the possible problem-settings selects one of them at random. This is followed by the unitary computation of the corresponding solution, which of course does not change the problem-setting selected by the initial measurement, and the final measurement of the solution, which does not change the state immediately before it.
@@ 3.1.1 Grover quantum search algorithm as a quantum superposition of causal loops
We called this quantum description ordinary because it is still the customary, causal, quantum description that cannot describe causal loops.
We observed that, involving only the measurements of commuting observables (the problem-setting and the solution), quantum algorithms are submitted to the same (classical) logic of the logical implication in question. Therefore, Grover algorithm should host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its optimal quantum speedup. Of course, this is not represented by its present ordinary quantum description, but we show that it is after completing this quantum description by time-symmetrizing it.
This time-symmetrization is, of course, submitted to the requirement of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As we shall see, the only possible way of obtaining this quantum superposition is by requiring that the initial and final measurements time-symmetrically contribute to the selection of the sorted out pair of one-to-one correlated measurement outcomes.
Things should go as follows. In each time symmetrization instance, the initial measurement selects one of the halves of the information that specifies the sorted out problem-setting (and thus also the corresponding solution) – instead of selecting a single problem-setting as customary, it selects a suitable quantum superposition of problem-settings. Its measurement outcome unitarily propagates forward in time becoming, immediately before the final measurement, a quantum superposition of tensor products, each one of the problem-settings selected by the initial measurement multiplying the corresponding solution. The final measurement of the solution completes the selection. Note that, encoding the tensor product of the sorted out problem-setting and the corresponding solution, the final measurement outcome is the same of the ordinary quantum description. Eventually, time-symmetrically, this measurement outcome propagates backward in time, by the inverse of the time-forward unitary evolution, until replacing the previous initial measurement outcome.
At this point, we should make two important observations:
(i) The backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the quantum algorithm. Disregarding the time-direction of causality, read from left to right, it is the unitary part of the ordinary quantum description of the quantum algorithm back again. Since this is the same in each time-symmetrization instance, also the quantum superposition of all instances is that unitary part.
Summing up, we can say that the time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered. What was incomplete is the physical part of the quantum description, that is, the description of the behavior of causality along the unitary part of the quantum process.
(ii) In each time-symmetrization instance, the forward and the backward in time propagations, together, constitute a causal loop that changes its own past (the outcome of the initial measurement). As we will see in the formalizations of subsection 3.2, this also changes the initial complete ignorance of the solution on the part of the problem-solver into knowledge of one of the possible halves of the information about it. There is a corresponding reduction of the computational complexity of the problem to be solved and the dimension of the Hilbert space in which the search for the solution takes place. The number of computation steps needed to solve the problem in an optimal quantum way is the number needed to solve in an optimal logical way the reduced computational complexity problem. We are of course dealing with one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
It is easy to show that the above time-symmetrization is the only possible way of generating the causal loops whose quantum superposition the quantum algorithm must host.
In fact, to have a causal loop, in the first place the selection of the information that specifies the sorted out pair of correlated measurement outcomes must properly share between the initial and final measurements with the selection performed by the initial measurement propagating forward in time until becoming the state immediately before the final measurement and that successively performed by the final measurement propagating backward in time until replacing the previous initial measurement outcome. Of course, this creates the forward and backward in time parts of a causal loop. Moreover, to (exactly) generate the causal loops logically implied by the sheer existence of the optimal quantum speedup, the sharing should be even.
Note that this is nothing but Laplace’s demon brought into the quantum world. Here, due to the possibility of having quantum state superpositions, there is more freedom. Indeed, it is a freedom exploited by the time-symmetrization process to create the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Eventually, let us observe that the quantum causal loops we are dealing with are generated by a mandatory time-symmetrization procedure – mandatory to generate the causal loops that must be hosted in the quantum algorithm. Therefore, as long as quantum mechanics is consistent, these causal loops should be exempt from the paradoxes that plague the causal loops of the usual kind. When needed to distinguish them from the latter kind, we will call them time-symmetric causal loops because generated by a time-symmetrization procedure.
3.1.2 Generalization
All the above has been developed for Grover algorithm. However, let us show how it generalizes. We should make two premises:
(i) All the results achieved until now have been obtained with no need of knowing the unitary evolution between the two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement – i.e. the quantum algorithm itself. We only needed to know that such a unitary evolution exists. This was the case until now, the unitary evolution being that of Grover algorithm (it will always be the case in the following).
(ii) Since the eigenvalue selected by the initial measurement remains unchanged throughout the unitary evolution in question, one is free to postpone that measurement at will. When performed simultaneously with the final measurement, the time-symmetrization we are dealing with must already have been put in place to represent the perfect (atemporal) symmetry between the simultaneous measurements of two maximally entangled, one-to-one correlated, observables. This is a second motivation for time-symmetrizing the ordinary quantum description of the quantum process in question, independent of the need of reproducing in the quantum algorithm the causal loops logically implied by the sheer existence of the quantum speedup. There is, of course, a redundancy between the two motivations. We produced both because together they say more.
Putting (i) and (ii) together, brings us well beyond the optimal Grover algorithm considered until now. In fact we obtain that:
The complete quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing it.
We call the above the causal loop rule of quantum correlation. Let us examine a few of its consequences.
In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal444 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting. . The time-symmetrization of its ordinary quantum description always yields that it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. In equivalent terms, there is a corresponding reduction of the computational complexity of the problem to be solved, and the dimension of the Hilbert space in which the search for the solution (or, identically, the one-to-one correlation) takes place. However, any of these facts just creates the room for a quantum speedup. To also have a quantum speedup, it takes that the problem-solver, or more generally the unitary evolution, exploits such a room. The room is fully exploited in the optimal quantum case, but it can also be not exploited at all or only partly exploited.
The latter is the case of the random quantum search algorithm $\left[11\right]$ that yields a suboptimal quantum speedup. Later, we will need to make reference to this quantum algorithm.
Of course, the quantum process may not even be a quantum algorithm. For example, quantum nonlocality is a process of this kind.
It in fact consists of a unitary evolution (the process of spatial separation between the two subsystems) between two one-to-one correlated measurement outcomes – those of measuring in the same basis the two maximally entangled observables. The process of spatial separation, of course, never changes the eigenvalue selected by any measurement. Here, each causal loop locally goes from one to the other measurement outcome via the time the two subsystems were not yet spatially separate where, always locally, it changes the state of both observables in a way that ensures the appropriate correlation between the two future measurement outcomes – see $\left[9\right]$. This is exactly the role that Einstein and the others ascribed to the hidden variables they envisaged.
Thus, the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and the others, in their (EPR) paper, ascribed to the quantum description of quantum nonlocality. Of course, in both cases what is incomplete is the ordinary quantum description of a quantum correlation.
By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[33\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[34\right]$. We do not understand why the argument concerning the incompleteness of the quantum description, central in that paper, has been removed by the scientific community. As is also well known, the question of the incompleteness of the quantum description has been at the center of the famous Einstein-Bohr dispute. Bohr, who was for the completeness, was eventually decreed the winner by the scientific community. Apparently that was once and for all.
According to the present line of thinking, also this position of science should be changed. To us, of course, Einstein should have been the winner. Another time, he would have put his finger on a very sensitive point in the quantum mechanical theory.
Eventually, as we shall see in Section 4., under a quantum cosmological interpretation of the Anthropic Principle, also the evolution of the universe toward the value of life turns out to be a unitary evolution between two one-to-one correlated measurement outcomes (the value of the fundamental constants since the beginning of time and the value of life today) that does not change the eigenvalue selected by the initial measurement (the value of the fundamental constants). As such, even if the search for the value of life by the universe’s evolution under the fundamental physical laws and constants were random, it would be endowed with quantum speedup and therefore teleological in character.
3.2 Formalization
We formalize the notion of the time-symmetric causal loops on the simplest, four drawer, instance of Grover quantum search algorithm. At the end of this subsection we will see how this immediately generalizes. The formalization consists of the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We call this extended description the ordinary quantum description, of the quantum algorithm: ordinary because it is still a causal description that cannot describe causal loops.